GLAS setup/observing recipe
NAOMI · WHT · Astronomy · ING

Draft 2007 Aug 30 - *** indicates query or unresolved issue

The recipes below supplement the NAOMI setup and observing recipes.

A separate page lists known GLAS-related problems, and things to fix/improve.

Safety

While the laser is powered on, nobody should enter the dome without wearing safety glasses.

The laser is mounted on the top-end ring of the WHT, feeding the beam-launch telescope behind the secondary mirror. The intense 5149-A light from the laser is an eye-safety hazard anywhere in the dome area, particularly if any part of the beam reflects (accidentally, or intentionally) off the inside surface of the dome.

Access to the dome is strictly controlled (by the laser safety duty officer) during observing.

Startup

To switch to LGS WFS mode:

To switch to NGS mode, follow a similar procedure, but select NAOMI (not ELECTRA) rather than GLAS, and select master instead of slave.

Note that on the LGS WFS, there are only 8*8 spots (and the bottom four are very faint), due to vignetting. Each spot is accompanied by a ~ 1%? ghost due to reflection from the back surface of the beam-splitter.

The laser is started up by the daytime engineering staff (ideally at least 6 hours before observing). The DEs' laser start-up and laser/BLT checks instructions can be found on the ops-team home page.

The laser traffic-control web page is usually brought up with the observing system.

To kill a window on aodisplay, you can no longer use right mouse button, 'kill window', like you could on solaris: 1.Type xkill in a terminal. 2.A message appears saying Select the window whose client you wish to kill with button 1...., and a skull and bones pointer appears. Click on the appropriate window with the left mouse button. (If you change you mind, then click on one of the other mouse buttons.)

Mirror-flattening

The DM is flattened by following the usual laser, white-light (NGS WFS) and simplex procedures. These ensure that a point source at infinity (simulated by the pinhole in the Nasmyth calibration unit) will yield a sharp image on the science detector.

At this stage, if required, the DM / NGS loop (i.e. just about everything apart from GLAS) can be tested by measuring offsets on the NGS WFS, closing the NGS WFS loop and checking what happens the pinhole is shifted 0.1 mm (the usual test).

For high-order correction with the LGS, it's necessary to measure offsets on the LGS (see below), *not* on the NGS.

LGS WFS

The LGS WFS sits behind the DM. Laser light reaches the LGS-WFS via the Laser Steering Mirror (LSM) and the Pockels cell shutter. For light from the laser calibration source to reach the LGS-WFS:

Measuring the WFS offsets

Once the DM has been flattened, the offsets on the LGS WFS can be measured by illuminating it with light from the LGS calibration source (pinhole, just in front of the NCU), which simulates a point source at the expected height of the laser:

Taking offsets on-sky (27/8/07):
To obtain a good AO correction, the best strategy was to take offsets
on-sky. To do this, (i) centre the LGS on the WFS using the BLT, and
(ii) use the LSM closed loop to take out any motion, then (iii)
measure the offsets. (The spots should be as well centred as possible
to maximise the dynamic range.)

At the moment, the LSM loop is only controllable by close the LGS
loop. Therefore Tim had to change the gains to 'freeze' the DM, so
that the offsets could be taken while the LGS loop was closed:

WFS SetGain 0.0002 (NB. Unilluminated segments stay fixed at 0.)
WFS SetLSMgain 5000 5000

He then set up plots of segment 1 (LSM) and a different segment (one
of DM segments) to check that only the LSM was moving, not the DM,
once the LGS loop was closed.

2.5 ms exposure time on WFS.

Check 'None' loaded for Offsets.
WFS ZeroWFSoffsets slave [he wasn't 100% sure if next command did this anyway]
WFS SetWFSoffsetsToCurrentCentroids slave
WFS SaveWFSoffsets slave 

If a large offset on one segment, then open the loop, maybe increase
to 5ms integration time, then close the loop and repeat taking the
offsets.

N.B. Tim originally tried using larger LSM gain and smaller DM gain,
but numerical errors caused problems.  Also, half way through the 100
frames taken for the offsets, he lowered the DM gain to 0.0001.

What should the gains be reset to?

[Aside: Currently the LSM is not offloading to the BLT.]

To load offsets (e.g. if you are testing the AO correction for offsets
taken on-sky vs off-sky), use TopGui. (There is also a command line
option, but Tim is not sure whether it works.)

GLAS control

Below is a snapshot of the GLAS control gui:

Click here for the web page on which the GLAS internal-camera images are displayed. Alternatively, type 'gracecamserver1' in a web browser (this currently does not work on aodisplay - try using whticsdisplay1).

GLAS light-path mimic

Some aspects of the GLAS LPM have been commissioned while others need to be completed (some require other hardware/software changes first). The current status is given below:
1.The laser ON/OFF status is not known by the observing system
yet. For the purposes of commissioning, the mimic now has the status
fixed to be ON all the time.  

2.The beam dump should show the correct
status. (However do not rely on this at this point.)  

3.The boresight
camera mirror should show the correct status, but it hasn't been
possible to thoroughly test this.  

4.The BLT x, y, focus values are
correctly displayed. (At a later date we may ask Craige to change the
precision, but it depends if we continue to use the x, y values or if
a different coordinate system is used.)  

5.The power values should be
ignored for now, as this is awaiting further engineering work.  

6.The
Laser calibration unit is deployed manually, therefore is fixed to be
'out' on the mimic.  

7.The Pockels cell and Image/Pupil View mirror
status are not known by the observing system.  

8.No LGS WFS info is
displayed yet.  

9.The NGS WFS status is correctly displayed.  

10.There
is no LSM information (I need to check the original specifications).

11.The tiptilt info and NGS loop status is correctly displayed.

12.The LGS loop status isn't being updated.  

13.In the NGS WFS, the
(i) camera focus, (ii) lenslet and (iii) filter mechs update okay. I
don't think the pickoff probe does, and I haven't tested the
foreoptics focus (although I suspect this will be okay).  
14.Most of
the realtime status (certainly the mirror flat and the DM gain) seemed
to be okay apart from the SG gain.
On Aug 27, the laser light status was correctly displayed up to the Pockels cell. Between the Pockels cell and the WFS the laser light was displayed in a warning colour because the mimic had a false Pockels cell status that did not match the requirements for on-sky operations.

On-sky laser checks

Tip-tilt correction

The scale is 0.25 arcse/pixel?

Open Jure's TT control page. To start the TT server, using sextans (with an obssys applied):

(This will be part of GlasRestart in future.)

Note that a log created by the TT control system can be found at tiptilt/2007xxyy. A new log is started every day at 12 UT. Severe problems are colour coded in red.

At the start of the night you should take bias frames (using the 'Take bias' button) for the TT configurations you plan to use.

Note that the tiptilt real time process is automatically started again if it stopped (e.g. with the 'kill' button on the GUI). If the controller is switched off, the RT process goes to sleep. It will recover when the controller is switched on again.

Below are screen dumps of the top part of the page in camera and plotter modes respectively:

Click on topgui NGS display to bring up tit-tilt display at centre of main WFS display (currently not very useful, intensity levels are wrong).

Observing

The acqtool has been calibrated (with dichroic 4). When the pickoff is at 0,0, the acqtool pixel is 286.0 283.5. The acqtool has successfully been used, but some of the OASIS acquisition features may not work if they require control of the AO loop.

To perform open loop tests, but with autoguiding, keep the TT loop closed, but with the gain set to zero. (Jure can give more details.)

End-of-night

Follow the 'turn laser off' instructions on the laser safety duty officer checklist.

Follow the usual end-of night procedures for NAOMI and the science instruments, and cover the GLAS optics.

Command glossary

Troubleshooting

TopGui very slow to respond.
Kill duplicate processes, restart topgui (type: ps -ef | grep TopGui; kill XXX; TopGuiGlas &). Try running oasisgui on lpss94 instead of sextans.
Beam dump can't be taken out while parked at zenith.
There could be a discrepancy with the TCS (if telescope in ENG mode?), meaning that the laser safety system won't permit the beam dump to be moved out.???
Acqtool can't move the pickoff probe.
Ensure that the NAOMIHOST environmental variable on taurus has been set to 'sextans.ing.iac.es'. Check that the EPMDataServer is running on sextans.
Loss of c40 communications
Type ¡Èc40CommunicationsRestart¡É as naomi@aocontrol1
Sequencer dies
Type: Sequencer2.py
TopGui won't come up in LGS mode
GlasRestart
Serious GLAS RTC failure
Rene's notes 26/8/07:
- kill TopGuiGlas process
- GlasRestart
- Look out for error messages of scripts that fail to load. Currently three
fail always (this will be fixed; is easy).
- To load the three failing scripts, enter the Python script names in
  sextans xterms:
     Sequencer2.py (this should be run from a separate terminal)
     SubscribeBridge.py (this and PublishBridge.py can be run from the same
       xterm)
     PublishBridge.py
- WFS SetGlasDefaults     to reload defaults
- TopGuiGlas   to run up TopGui
- on CPU page, close the strain gauge loop
- on CPU page, start LGS WFS
- If there is a problem with the framing it may 
  be necessary to re-init the 
  LGS WFS on the SDSU page. However do not do this by default.
- on mirror control page, LOAD last flat
- on real time page, LOAD NAOMI reconstructor
- on WFS page, load offsets
- on SDSU page, set integration time
- on an Xterm type: CameraSwitch.py and in the box that appears click
'slave'.
Check that slaveSynch is set as the camera. 
Then exit that box.
- on TopGuiGlas activate continuous readout and adjust display control if
necessary.

Technical information

Introductions to GLAS can be found in the ING newsletter Dec 2005 and on the GLAS instrument-development page.

The main components of the system are shown below (figure from the newsletter article):

The GLAS cradle (top-end ring) and the layout of the optics on the GRACE optical bench are shown below (also taken from the newsletter article):

Individual components in the GLAS launch system:

The beam is 35 cm wide when it leaves the BLT, and focuses to a spot 10 cm wide at altitude 15 km (i.e. subtended angle ~ 1 arcsec). The laser pulses at 5000 HZ?, and each pulse lasts 400 nsec?, total power ~ 17 - 19 W, yielding an artificial star with equivalent mag ~ 9.5?

Individual components in the GLAS LGS WFS light path:

Useful numbers:

NAOMI · WHT · Astronomy · ING

ING Logo

Last Updated: 2007 Aug 26
Chris Benn crb@ing.iac.es, Sam Rix srix@ing.iac.es Rene Rutten rgmr@ing.iac.es